# Neural Force Field: Few-shot Learning of Generalized Physical Reasoning

### v4

@@ Appendix E Planning time on N-body
Appendix F Detailed planning results
We present the detailed planning results of different models across trials in Tab.˜A5.
@@ Appendix F Detailed planning results
We test the performance of GPT-4o for both within-scenario planning and cross-scenario planning. Detailed results are presented in Tab.˜A6. Within-scenario games contains the same 10 games used in the prompt in Round 1. So the language model can memorize corresponding solutions provided in the prompt to achieve good performance. Cross-scenario games contains the other 30 game settings, on which the performance of GPT-4o is better than other baselines (Random, IN, SlotFormer) in Tab.˜A5. The improvement of performance through refinement is witnessed in both within-scenario and cross-scenario games.
@@
Appendix G More visualization
We provide additional visualizations of the learned forces along trajectories in I-PHYRE, as shown in Fig.˜A1, Fig.˜A2, and Fig.˜A3. Further prediction results on the N-body dataset are presented in Fig.˜A4. Additionally, examples of failure cases in NFF predictions are shown in Fig.˜A6. We also visualize the failure encoding of slot-attention in SlotFormer in Fig.˜A8.

### v5

@@ Appendix F Planning time on N-body
Appendix G Detailed planning results
We present the detailed planning results of different models across trials in Table A4.
@@ Appendix G Detailed planning results
We test the performance of GPT-4o for both within-scenario planning and cross-scenario planning. Detailed results are presented in Table A5. Within-scenario games contains the same 10 games used in the prompt in Round 1. So the language model can memorize corresponding solutions provided in the prompt to achieve good performance. Cross-scenario games contains the other 30 game settings, on which the performance of GPT-4o is better than other baselines (Random, IN, SlotFormer) in Table A4. The improvement of performance through refinement is witnessed in both within-scenario and cross-scenario games.
@@
Appendix H More visualization
We provide additional visualizations of the learned forces along trajectories in I-PHYRE, as shown in Figure A1, Figure A2, and Figure A3. Further prediction results on the N-body dataset are presented in Figure A4. Additionally, examples of failure cases in NFF predictions are shown in Figure A6. We also visualize the failure encoding of slot-attention in SlotFormer in Figure A8.